# Database for the meta-analysis of the social cost of carbon (v2025.1)

### v3

@@ Abstract
A new version of the database for the meta-analysis of estimates of the social cost of carbon is presented. New records were added, and new fields on gender and stochasticity.
@@ 1 Introduction
Previous meta-analyses of the social cost of carbon were published by Tol (2005b), van den Bergh and Botzen (2014), Havránek et al. (2015), Wang et al. (2019) and Moore et al. (2024). My first attempt was updated in Tol (2008, 2009, 2010b, 2011, 2013b, 2014, 2015, 2018, 2019a, 2020, 2023a, 2023b), Anthoff and Tol (2022) and Tol (2025). Many of the updates involve newer data only. The exceptions are Tol (2008) and Anthoff and Tol (2022), who focus on the tail of the distribution, Tol (2018) and Tol (2025) who try to detect publication bias (the latter successfully), and Tol (2008) and Tol (2023b), who study trends. Tol and Tol (2023) is about software for online visualization of meta-analysis.
The new version of the database includes new records and a few new fields. I shifted the base year to 2025, and now use 2024 price levels. I added gender [^1] Gender is the gender recorded at the time of publication. and whether the model is stochastic—the first such paper is Crost and Traeger (2013)—or solved analytically—Nordhaus (1980) used an analytical approximation of a steady-state economy and climate, but analytical integrated assessment models only took off following Golosov et al. (2014).
@@ 2.1 Literature
The current version of the database includes 446 papers, spanning the period from 1980 to 2024. These papers were collected over two decades. Only two papers are missing, the other 444 are in the author’s library, mostly in electronic form but also as hard copy. Papers were collected by systematic literature search, personal acquaintance, and exploration of the citations of previously considered papers. The latest database also includes many papers found by Moore et al. (2024) but hitherto overlooked by me.
@@
The papers included in the current version of the meta-analysis are: Ackerman and Finlayson (2006), Ackerman and Munitz (2012, 2016), Ackerman and Stanton (2012), Ackerman et al. (2013), Adler et al. (2017), Agliardi and Xepapadeas (2022), Allen (2016), Anderson et al. (2014), Anthoff (2007, 2009), Anthoff and Emmerling (2019), Anthoff and Tol (2009, 2010, 2011, 2013, 2014, 2022), Anthoff et al. (2009a, b, c, 2011a, 2011b), Asplund (2017, 2019), Audus (1998), Ayres and Walter (1991), Azar (1994), Azar et al. (2023), Azar and Sterner (1996), Balmford et al. (2023), Barnett et al. (2020), Bastien-Olvera and Moore (2021), Barrage (2014, 2018, 2020b, 2020a), Barrage and Nordhaus (2023, 2024), Bauer and Rudebusch (2023), Belfiori and Rezai (2024), von Below (2014), Berger et al. (2017), Bertram et al. (2024), Bherwani et al. (2019, 2024), van den Bijgaart et al. (2013, 2016), Bond (2024), Botzen and van den Bergh (2012), Brander et al. (2010), Braun et al. (2024), Brausmann and Bretschger (2024), van den Bremer and van der Ploeg (2021), Bressler (2021), Bretschger and Karydas (2018), Bretschger and Pattakou (2019), Brock and Xepapadeas (2017, 2019), Budolfson and Dennig (2020), Budolfson et al. (2017, 2019), Caesary et al. (2023), Cai and Lontzek (2019), Cai et al. (2012, 2015, 2016), Calvas et al. (2024), Ceronsky et al. (2005, 2011), Chung et al. (2024), Clarkson and Deyes (2002), Cline (1992, 1997, 2004), Coleman et al. (2021), Coppola et al. (2024), Crost and Traeger (2013, 2014), Czyz and Safarzyńska (2023), Dangl and Wirl (2007), Daniel et al. (2019), Dayaratna and McKitrick (2023), Dayaratna et al. (2017, 2020), Dennig (2014), Dennig et al. (2015), Dietz (2011), Dietz and Koninx (2022), Dietz et al. (2021a, b), Dietz and Stern (2015), Dietz and Venmans (2019), Dong et al. (2024), Downing et al. (1996, 2005), Drupp and Hänsel (2021), Ekholm (2018), EPA (2023), EPA and NHTSA (2009), Espagne et al. (2018), Eyre et al. (1999), Fankhauser and Pearce (1993), Fankhauser (1994, 1995), Farooq and Cora (2024), Faulwasser et al. (2018b, a), Fiestas-Chevez et al. (2024), Fillon et al. (2023), Foley et al. (2013), Freeman and Groom (2015), Freeman et al. (2015), Freeman and Groom (2016), Fu et al. (2023), Gerlagh (2023), Gerlagh et al. (2020), Gerlagh and Liski (2017, 2018), Gerlagh et al. (2023), Gillingham et al. (2018), Glotter et al. (2014), Golosov et al. (2014), Golub and Brody (2017), González-Eguino and Neumann (2016), Gössling and Humpe (2024), Goulder and Mathai (2000), Greenstone et al. (2013), Griffiths et al. (2012), Groom et al. (2007), Groom and Venmans (2023), Groshans et al. (2019), Gschnaller (2020), Guivarch and Pottier (2018), Guo et al. (2006), Hafeez et al. (2017), Haites (1998), Hambel et al. (2021b, a, 2024), Hanley and Tinch (2004), Hänsel and Quaas (2018), Hänsel et al. (2020, 2021), Haraden (1992, 1993), Hasan et al. (2024), Hassler et al. (2016), Hastunç et al. (2024), Hatase and Managi (2015), Heal and Millner (2014), Heinzow and Tol (2003), Hepburn et al. (2009), Hillebrand and Hillebrand (2019, 2023), Hoffman et al. (2024), Hohmeyer (1996, 2002, 2004), Hohmeyer and Gaertner (1992), Hong et al. (2023), Hope (2005a, b, c, 2006a, 2006b, 2006c, 2008a, 2008b, 2009, 2011b, 2011a, 2013a, 2013b, 2020), Hope and Hope (2013), Hope and Maul (1996), Howard and Sterner (2017), Howard and Sylvan (2020), Howarth (1998), Howarth et al. (2014), Hwang et al. (2013, 2016, 2017), Hwang (2017), Hwang et al. (2019), Interagency Working Group on the Social Cost of Carbon (2010, 2013, 2015, 2021), Ibrahim et al. (2024), Ikefuji et al. (2021), Iverson (2012), Iverson et al. (2015), Iverson and Karp (2021), Jaakkola and Millner (2022), Jaakkola and van der Ploeg (2019), Hoffman et al. (2024), Al-Jabir and Isaifan (2024), Jensen and Traeger (2013, 2014, 2021), Jiang et al. (2024), Jin et al. (2024), Johnson and Hope (2012), Kalkuhl and Wenz (2020), Karydas and Xepapadeas (2019), Kaushal and Navrud (2023), Kellett et al. (2019), Khabarov et al. (2022), Kelly and Kolstad (2001), Kemfert and Schill (2010), Kessler (2017), Kikstra et al. (2021), Knoke et al. (2023), Koch and Leimbach (2023), Kon and Caner (2024), Kopp et al. (2012), Kotchen (2018), Kotlikoff et al. (2021), Krewitt and Schlomann (2006), Kulkarni et al. (2024), Lemoine (2015, 2021), Lemoine and Traeger (2014, 2016a, 2016b), Li et al. (2016), Link and Tol (2004), Lintunen and Rautiainen (2021), Liu et al. (2022), Lontzek et al. (2015), Loube (2024), Lucchesi et al. (2024), Lupi and Marsiglio (2021), Maddison (1994, 1995), Manne (2004), Mardones (2024), Marten (2011, 2014), Marten et al. (2015), Marten and Newbold (2012a, b, 2013), Mastandrea and Schneider (2001), Mendelsohn (2003, 2004, 2005), Meng et al. (2024), van der Mensbrugghe (2023), Mikhailova et al. (2019, 2024b, 2024a), MnPUC (2018), Molocchi and Mela (2024), Moore and Diaz (2015), Moore et al. (2017), Moyer et al. (2014), Muangjai et al. (2024), Naeini et al. (2020), Narita et al. (2009, 2010a, 2010b), Newbold et al. (2010, 2013), Newbold and Marten (2014), Newell and Pizer (2003, 2004), Newell et al. (2022), Niu and Zou (2024), Nordhaus (1980, 1982, 1989, 1990a, 1990b, 1991a, 1991b, 1991c, 1992, 1993a, 1993b, 1993c, 1994b, 1994a, 1997, 2001, 2007b, 2007c, 2007a, 2008, 2009, 2010, 2011, 2013a, 2013b, 2014, 2015, 2017b, 2017a, 2018b, 2018a, 2019b, 2019a), Nordhaus and Boyer (2000), Nordhaus and Popp (1997), Nordhaus and Sztorc (2013), Nordhaus and Yang (1996), NYSERDA and RFF (2021), Okullo (2020), Olijslagers (2022), Olijslagers et al. (2023), Olijslagers and van Wijnbergen (2024), Ortiz et al. (2011), Otto et al. (2013), Parry (1993), Pearce (2003, 2005), Peck and Teisberg (1992, 1993, 1994, 1995, 1996), Penner et al. (1992), Perrissin Fabert et al. (2012), Pezzey and Burke (2014), Pindyck (2017, 2019), Piontek et al. (2019), Pizer (1999, 2002), Plambeck and Hope (1996), van der Ploeg (2014, 2015), der Ploeg and Rezai (2016), van der Ploeg and Rezai (2017), van der Ploeg and Rezai (2018), van der Ploeg and Rezai (2019b, a), van der Ploeg et al. (2012), van der Ploeg (2018), van der Ploeg and Rezai (2021), van der Ploeg (2021), van der Ploeg and de Zeeuw (2013), de Zeeuw and van der Ploeg (2014), van der Ploeg and de Zeeuw (2016, 2017, 2019), Poelhekke (2019), Popp (2004), Portland and South Portland (2021), Pottier et al. (2015), Prest and Stock (2023), Prest et al. (2024), Price et al. (2007), Pulhin et al. (2024), Pycroft et al. (2011, 2014), Quiggin (2018), Rautiainen and Lintunen (2017), Reilly and Richards (1993), Rennert and Kingdon (2019), Rennert et al. (2021, 2022), Rezai (2010), Rezai and van der Ploeg (2015, 2016, 2017b, 2017c, 2017a), Rezai et al. (2012, 2020), Ricke et al. (2018), Rickels et al. (2023, 2024), Rose et al. (2017), Roughgarden and Schneider (1999), Rudik (2020), Russell et al. (2022), Di Russo et al. (2024), Safarzyńska and van den Bergh (2022), Schauer (1995), Schmidt et al. (2013), Schultes et al. (2021), Scovronick et al. (2017, 2020), Shafiei-Alavijeh et al. (2024), Sen et al. (2024), Shindell (2015), Shindell et al. (2017), Smith et al. (2023), Sohn (2019), Sohngen (2010), Stern et al. (2006), Stern and Taylor (2007), Striepe et al. (2024), Su et al. (2017, 2024), Taconet et al. (2021), Tao et al. (2024), Tian et al. (2019), Tibebu et al. (2021, 2024), Tol (1999, 2003, 2005a, 2005c, 2010a, 2010c, 2012b, 2013a, 2013c, 2019b), Tol and Downing (2000, 2001), Tol (2020, 2024), Traeger (2014, 2015, 2023), Umweltbundesamt (2007, 2012, 2019, 2020), Uzawa (2003), Wahba and Hope (2006), Waldhoff et al. (2011, 2014), Wang et al. (2022, 2024), Watkiss (2005), Watkiss and Downing (2008), Weitzman (2013), Weller et al. (2015), van der Wijst et al. (2021), van der Wijst et al. (2023), Wong et al. (2015), Wu et al. (2024), Yang et al. (2018), Yilma and Yitay (2024), Yohe and Hope (2013), Yoo and Mendelsohn (2018, 2021), Zhao et al. (2023), Zhen and Tian (2020) and Zhen et al. (2018).
@@ 2.2 Main database
Results and assumptions were coded for each of the 446 papers. In most cases, tables and text were used. Some papers present results only graphically, in which case authors were contacted by email. If there was no response, graphical results were digitized with the grabit function in Matlab.
The main database is in Excel. It consists of the following fields, for each of the 14,152 estimates of the social cost of carbon:
@@
    The social cost of carbon for 2010 in 2010 US dollar per metric tonne of carbon.
@@
    New The social cost of carbon for 2025 in 2024 US dollar per metric tonne of carbon.
@@
    Reassessed Functional form of the level damage function.
@@
    newAssumed growth impact of 1.0 warming.
@@
    new Functional form of the growth damage function.
@@
    New Stochastic, a dummy variable whether the model used is stochastic (1) or not.
@@
    New Analytic, a dummy variable whether the model is solved analytically (1) or numerically (0).
@@
    Reassessed Dynamic, a variable indicating whether vulnerability to climate change falls with development (1), increases (-1), or does not change (0).
@@
A supplementary database records the social cost of carbon for different years of emission. This is used to calculate the population average growth rate, which is 1.97% per year in the current database. This growth rate is used to shift the social cost of carbon for the year of reporting to 2025.

### v4

# Database for the meta-analysis of the social cost of carbon (v2026.1)

@@ Abstract
A new version of the database for the meta-analysis of estimates of the social cost of carbon is presented. New records were added.
@@ 1 Introduction
Previous meta-analyses of the social cost of carbon were published by Tol (2005b), van den Bergh and Botzen (2014), Havránek et al. (2015), Wang et al. (2019) and Moore et al. (2024). My first attempt was updated in Tol (2008, 2009, 2010b, 2011, 2013b, 2014, 2015, 2018, 2019a, 2020, 2023a, 2023b), Anthoff and Tol (2022), Tol (2025b) and Tol (2025a). Many of the updates involve newer data only. The exceptions are Tol (2008) and Anthoff and Tol (2022), who focus on the tail of the distribution, Tol (2018) and Tol (2025a) who try to detect publication bias (the latter successfully), and Tol (2008) and Tol (2023b), who study trends. Tol and Tol (2023) is about software for online visualization of meta-analysis.
The new version of the database includes new records.
@@ 2.1 Literature
The current version of the database includes 528 papers, published in 156 journals, spanning the period from 1980 to 2025. These papers were collected over two decades. Only two papers are missing, the others are in the author’s library, mostly in electronic form some only in hard copy. Papers were collected by systematic literature search, personal acquaintance, and exploration of the citations of previously considered papers.
This version of the database adds 60 papers published in 2025: Akdağ (2025), Aksoy and Bayram Arlı (2025), Arab et al. (2025), Ashenfarb and Kotchen (2025), Behzadi and Arani (2025), Bhattacharya et al. (2025), Bhave et al. (2024), Brooks et al. (2025), Caesary et al. (2025), Castellini et al. (2025), Chapin et al. (2025), Chen et al. (2025a, b), Cheu and Gammans (2025), Cho et al. (2025), Coppens et al. (2025), Cunanan et al. (2025), da Costa et al. (2025), Deng et al. (2025), Eduardo et al. (2025), Estrada et al. (2025a, b), Fell et al. (2025), Feng (2025), Folini et al. (2025), García et al. (2025), Islam et al. (2025), Kashanian et al. (2025), Khan et al. (2025), Kerns et al. (2025), Kotchen (2025), Laamouri and Khattabi (2025), Landry (2025), Lima et al. (2025), Lubis et al. (2025), Manan et al. (2025), Mohammadyari et al. (2025), Naef et al. (2025), Neal et al. (2025), Ovaere and Proost (2025), Piazza et al. (2025), Van Der Ploeg (2025), Pranita and Sarjana (2025), Rao et al. (2025), Rose et al. (2025), Schaumann and de Asenjo (2025), Serfas (2025), Springmann et al. (2025), Sreelekshmi et al. (2025), Wang et al. (2025), Wu (2025), Wu et al. (2025a, b), Xiao et al. (2025), Xu et al. (2025), Xuan et al. (2025), Yu et al. (2025), Zhang et al. (2025), Zhu et al. (2025).
I also added 21 previously overlooked papers: Al-Amin et al. (2015), Aleshina et al. (2024), Bilal and Känzig (2024), Government of Canada (2022), Deng et al. (2024), Sylvan (2017b), Eteriki et al. (2023), GAO (2020), Howard and Sylvan (2015a, b), Landry (2021), Li et al. (2025), Rasiah et al. (2016), Raynaud (2014), Rose et al. (2022), Sarkar et al. (2019), Shelanski and Obstfeld (2015), Siagian et al. (2024), Sylvan (2017a), Umweltbundesamt (2024), Wong et al. (2016).
@@
These paper add to 446 papers included in the previous version of the meta-analysis: Ackerman and Finlayson (2006), Ackerman and Munitz (2012, 2016), Ackerman and Stanton (2012), Ackerman et al. (2013), Adler et al. (2017), Agliardi and Xepapadeas (2022), Allen (2016), Anderson et al. (2014), Anthoff (2007, 2009), Anthoff and Emmerling (2019), Anthoff and Tol (2009, 2010, 2011, 2013, 2014, 2022), Anthoff et al. (2009a, b, c, 2011a, 2011b), Asplund (2017, 2019), Audus (1998), Ayres and Walter (1991), Azar (1994), Azar et al. (2023), Azar and Sterner (1996), Balmford et al. (2023), Barnett et al. (2020), Bastien-Olvera and Moore (2021), Barrage (2014, 2018, 2020b, 2020a), Barrage and Nordhaus (2023, 2024), Bauer and Rudebusch (2023), Belfiori and Rezai (2024), von Below (2014), Berger et al. (2017), Bertram et al. (2024), Bherwani et al. (2019, 2024), van den Bijgaart et al. (2013, 2016), Bond (2024), Botzen and van den Bergh (2012), Brander et al. (2010), Braun et al. (2024), Brausmann and Bretschger (2024), van den Bremer and van der Ploeg (2021), Bressler (2021), Bretschger and Karydas (2018), Bretschger and Pattakou (2019), Brock and Xepapadeas (2017, 2019), Budolfson and Dennig (2020), Budolfson et al. (2017, 2019), Caesary et al. (2023), Cai and Lontzek (2019), Cai et al. (2012, 2015, 2016), Calvas et al. (2024), Ceronsky et al. (2005, 2011), Chung et al. (2024), Clarkson and Deyes (2002), Cline (1992, 1997, 2004), Coleman et al. (2021), Coppola et al. (2024), Crost and Traeger (2013, 2014), Czyz and Safarzyńska (2023), Dangl and Wirl (2007), Daniel et al. (2019), Dayaratna and McKitrick (2023), Dayaratna et al. (2017, 2020), Dennig (2014), Dennig et al. (2015), Dietz (2011), Dietz and Koninx (2022), Dietz et al. (2021a, b), Dietz and Stern (2015), Dietz and Venmans (2019), Dong et al. (2024), Downing et al. (1996, 2005), Drupp and Hänsel (2021), Ekholm (2018), EPA (2023), EPA and NHTSA (2009), Espagne et al. (2018), Eyre et al. (1999), Fankhauser and Pearce (1993), Fankhauser (1994, 1995), Farooq and Cora (2024), Faulwasser et al. (2018b, a), Fiestas-Chevez et al. (2024), Fillon et al. (2023), Foley et al. (2013), Freeman and Groom (2015), Freeman et al. (2015), Freeman and Groom (2016), Fu et al. (2023), Gerlagh (2023), Gerlagh et al. (2020), Gerlagh and Liski (2017, 2018), Gerlagh et al. (2023), Gillingham et al. (2018), Glotter et al. (2014), Golosov et al. (2014), Golub and Brody (2017), González-Eguino and Neumann (2016), Gössling and Humpe (2024), Goulder and Mathai (2000), Greenstone et al. (2013), Griffiths et al. (2012), Groom et al. (2007), Groom and Venmans (2023), Groshans et al. (2019), Gschnaller (2020), Guivarch and Pottier (2018), Guo et al. (2006), Hafeez et al. (2017), Haites (1998), Hambel et al. (2021b, a, 2024), Hanley and Tinch (2004), Hänsel and Quaas (2018), Hänsel et al. (2020, 2021), Haraden (1992, 1993), Hasan et al. (2024), Hassler et al. (2016), Hastunç et al. (2025), Hatase and Managi (2015), Heal and Millner (2014), Heinzow and Tol (2003), Hepburn et al. (2009), Hillebrand and Hillebrand (2019, 2023), Hoffman et al. (2024), Hohmeyer (1996, 2002, 2004), Hohmeyer and Gaertner (1992), Hong et al. (2023), Hope (2005a, b, c, 2006a, 2006b, 2006c, 2008a, 2008b, 2009, 2011b, 2011a, 2013a, 2013b, 2020), Hope and Hope (2013), Hope and Maul (1996), Howard and Sterner (2017), Howard and Sylvan (2020), Howarth (1998), Howarth et al. (2014), Hwang et al. (2013, 2016, 2017), Hwang (2017), Hwang et al. (2019), Interagency Working Group on the Social Cost of Carbon (2010, 2013, 2015, 2021), Ibrahim et al. (2024), Ikefuji et al. (2021), Iverson (2012), Iverson et al. (2015), Iverson and Karp (2021), Jaakkola and Millner (2022), Jaakkola and van der Ploeg (2019), Hoffman et al. (2024), Al-Jabir and Isaifan (2024), Jensen and Traeger (2013, 2014, 2021), Jiang et al. (2024), Jin et al. (2024), Johnson and Hope (2012), Kalkuhl and Wenz (2020), Karydas and Xepapadeas (2019), Kaushal and Navrud (2023), Kellett et al. (2019), Khabarov et al. (2022), Kelly and Kolstad (2001), Kemfert and Schill (2010), Kessler (2017), Kikstra et al. (2021), Knoke et al. (2023), Koch and Leimbach (2023), Kon and Caner (2024), Kopp et al. (2012), Kotchen (2018), Kotlikoff et al. (2021), Krewitt and Schlomann (2006), Kulkarni et al. (2024), Lemoine (2015, 2021), Lemoine and Traeger (2014, 2016a, 2016b), Li et al. (2016), Link and Tol (2004), Lintunen and Rautiainen (2021), Liu et al. (2022), Lontzek et al. (2015), Loube (2024), Lucchesi et al. (2024), Lupi and Marsiglio (2021), Maddison (1994, 1995), Manne (2004), Mardones (2024), Marten (2011, 2014), Marten et al. (2015), Marten and Newbold (2012a, b, 2013), Mastandrea and Schneider (2001), Mendelsohn (2003, 2004, 2005), Meng et al. (2024), van der Mensbrugghe (2023), Mikhailova et al. (2019, 2024b, 2024a), MnPUC (2018), Molocchi and Mela (2024), Moore and Diaz (2015), Moore et al. (2017), Moyer et al. (2014), Muangjai et al. (2024), Naeini et al. (2020), Narita et al. (2009, 2010a, 2010b), Newbold et al. (2010, 2013), Newbold and Marten (2014), Newell and Pizer (2003, 2004), Newell et al. (2022), Niu and Zou (2024), Nordhaus (1980, 1982, 1989, 1990a, 1990b, 1991a, 1991b, 1991c, 1992, 1993a, 1993b, 1993c, 1994b, 1994a, 1997, 2001, 2007b, 2007c, 2007a, 2008, 2009, 2010, 2011, 2013a, 2013b, 2014, 2015, 2017b, 2017a, 2018b, 2018a, 2019b, 2019a), Nordhaus and Boyer (2000), Nordhaus and Popp (1997), Nordhaus and Sztorc (2013), Nordhaus and Yang (1996), NYSERDA and RFF (2021), Okullo (2020), Olijslagers (2022), Olijslagers et al. (2023), Olijslagers and van Wijnbergen (2024), Ortiz et al. (2011), Otto et al. (2013), Parry (1993), Pearce (2003, 2005), Peck and Teisberg (1992, 1993, 1994, 1995, 1996), Penner et al. (1992), Perrissin Fabert et al. (2012), Pezzey and Burke (2014), Pindyck (2017, 2019), Piontek et al. (2019), Pizer (1999, 2002), Plambeck and Hope (1996), van der Ploeg (2014, 2015), der Ploeg and Rezai (2016), van der Ploeg and Rezai (2017), van der Ploeg and Rezai (2018), van der Ploeg and Rezai (2019b, a), van der Ploeg et al. (2012), van der Ploeg (2018), van der Ploeg and Rezai (2021), van der Ploeg (2021), van der Ploeg and de Zeeuw (2013), de Zeeuw and van der Ploeg (2014), van der Ploeg and de Zeeuw (2016, 2017, 2019), Poelhekke (2019), Popp (2004), Portland and South Portland (2021), Pottier et al. (2015), Prest and Stock (2023), Prest et al. (2024), Price et al. (2007), Pulhin et al. (2024), Pycroft et al. (2011, 2014), Quiggin (2018), Rautiainen and Lintunen (2017), Reilly and Richards (1993), Rennert and Kingdon (2019), Rennert et al. (2021, 2022), Rezai (2010), Rezai and van der Ploeg (2015, 2016, 2017b, 2017c, 2017a), Rezai et al. (2012, 2020), Ricke et al. (2018), Rickels et al. (2023, 2024), Rose et al. (2017), Roughgarden and Schneider (1999), Rudik (2020), Russell et al. (2022), Di Russo et al. (2025), Safarzyńska and van den Bergh (2022), Schauer (1995), Schmidt et al. (2013), Schultes et al. (2021), Scovronick et al. (2017, 2020), Shafiei-Alavijeh et al. (2024), Sen et al. (2024), Shindell (2015), Shindell et al. (2017), Smith et al. (2023), Sohn (2019), Sohngen (2010), Stern et al. (2006), Stern and Taylor (2007), Striepe et al. (2024), Su et al. (2017, 2024), Taconet et al. (2021), Tao et al. (2024), Tian et al. (2019), Tibebu et al. (2021, 2024), Tol (1999, 2003, 2005a, 2005c, 2010a, 2010c, 2012b, 2013a, 2013c, 2019b), Tol and Downing (2000, 2001), Tol (2020, 2024), Traeger (2014, 2015, 2023), Umweltbundesamt (2007, 2012, 2019, 2020), Uzawa (2003), Wahba and Hope (2006), Waldhoff et al. (2011, 2014), Wang et al. (2022, 2024), Watkiss (2005), Watkiss and Downing (2008), Weitzman (2013), Weller et al. (2015), van der Wijst et al. (2021), van der Wijst et al. (2023), Wong et al. (2015), Wu et al. (2024), Yang et al. (2018), Yilma and Yitay (2024), Yohe and Hope (2013), Yoo and Mendelsohn (2018, 2021), Zhao et al. (2023), Zhen and Tian (2020) and Zhen et al. (2018).
@@ 2.2 Main database
Results and assumptions were coded for each of the 528 papers. In most cases, tables and text were used. Some papers present results only graphically, in which case authors were contacted by email. If there was no response—in one case, the authors responded that their hard disk had crashed—graphical results were digitized with the grabit function in Matlab.
The main database is in Excel. It consists of the following fields, for each of the 14,884 estimates of the social cost of carbon:
@@
    Legacy The social cost of carbon for 2010 in 2010 US dollar per metric tonne of carbon.
@@
    The social cost of carbon for 2025 in 2024 US dollar per metric tonne of carbon.
@@
    Functional form of the level damage function.
@@
    Assumed growth impact of 1.0 warming.
@@
    Functional form of the growth damage function.
@@
    Stochastic, a dummy variable whether the model used is stochastic (1) or not.
@@
    Analytic, a dummy variable whether the model is solved analytically (1) or numerically (0).
@@
    Dynamic, a variable indicating whether vulnerability to climate change falls with development (1), increases (-1), or does not change (0).
@@
A supplementary database records the social cost of carbon for different years of emission. This is used to calculate the population average growth rate, which is 2.07% per year in the current database. This growth rate is used to shift the social cost of carbon for the year of reporting to 2025.